OpenStudy (anonymous):

The table below shows the values of a linear function. What will be the value of the function when x equals 7?

OpenStudy (anonymous):



OpenStudy (anonymous):

A -41 B -33 C -45

OpenStudy (anonymous):

@petiteme

OpenStudy (anonymous):

is it A

OpenStudy (petiteme):

Since it's a linear function, you can find the function using y=mx+b where m is the slope and b is the y-intercept. Then you plug in for x.

OpenStudy (anonymous):

then whats my answer?

OpenStudy (anonymous):

its either a or c

OpenStudy (anonymous):

m=(y-y1)/(x-x1) m=(-21-(-17))/(2-1) m=(-21+17)/(2-1) m=(-4)/(1) m=-4 (y-y1)=m(x-x1) (y-(-17))=m(x-1) y+17=-4(x-1) y+17=-4x+4 y=-4x-13

OpenStudy (anonymous):

now, plug in 7 for x, and solve for y